OpenStudy (anonymous):

8 5/9 + 3 5/6 + 6 4/9

OpenStudy (anonymous):

Someone Please help I have a few more hours!!!!!!!

OpenStudy (timaashorty):

ignor the whole numbers for now ( 8, 3, & 6) We're focusing on the fractions. 5/9 + 5/6 + 5/9 In order to add, we have to convert all of them with one same denominator. By doing that, we can multiply both denominators. (6*9= 4) now time to convert. 5 *6 30 - = - 9 * 6 54 5*9 45 - = - 6 *9 54 5*6 30 - = - 9 *6 54 now add all together. 30/54 + 45/54 + 30/54= 1 17/18 now add the whole numbers. 8+3 + 6 = 17 + 1 17/18 = 18 17/18 Therefore your final answer is \[18 \frac{ 17}{ 18 }\]

OpenStudy (timaashorty):

@OwlGirl212

OpenStudy (timaashorty):

if you're in a rush or something. I suggest you to use http://www.wolframalpha.com . Just submit you're problem and you'll get the answer you need.

OpenStudy (timaashorty):

Oops your final answer is 18 5/6* my bad \[18 \frac{ 5 }{ 6 }\]